"A man invested a total of $22,000 in any two investments. He made a profit of 7% on the first investment and 2% on the second investment. If his total profit was $920, what was the amount of each investment?"
Is there a formula for this type of problem.

Answer by guapa(62)

Yes, there is.
Profit earned from 7% investment + Profit earned from 2% investment = $920
Let x represent the amount invested at 7%. Then 22.000-x represents the amount invested at 2%.
Now you can set up the equation:
7%(x)+2%(22.000-x) = $920
0,07x + 0,02(22.000-x)=$920
0,07x + 440 - 0,02x = 920
0,05x = 480
x = 9600 (amount invested at 7%)
22,000 - x = 12,400 (amount invested at 2%)
Check it!
(i=Prt)